A company produces two competing products whose demand schedules are
q1 = 219 - 1.8p1 + 0.5p2
q2 = 303 - 2.1p2 + 0.8p1
What price should it charge in the two markets to maximize total sales revenue?
1
Expert's answer
2020-04-15T09:47:51-0400

"TR=TR1+TR2=p1q1+p2q2;"

"p1\u00d7(219-1.8p1+0.5p2)+p2\u00d7(303-2.1p2+0.8p1);"

"=219p1-1.8p1^2+0.5p1p2+303p2-2.1p2^2+0.8p1p2;"

"1.3p1p2+219p1+303p2-1.8p1^2-2.1p2^2;"

"\frac {dTR}{dp1}=219-3.6p1+1.3p2;"

"\frac {dTR}{dp2}=303-4.2p2+1.3p1;"

Multiplying by 42

"9 198 - 151.2p1+54.6p2=0";

Multiplying by 13

"3939+16.9p1-54.6p2=0;"

"13137-134.3p1=0;"

p"1=98;"

"303-2.1p2+0.8\u00d798; 2.1p2=303+78.4; 2.1p2=381.4; p2= 182;"

Answer : p1= $98 and p2= $ 182.